# Jamais Vu: Exposing the Generalization Gap in Supervised Semantic Correspondence

### v1

@@ Appendix A Additional Implementation Details
We base our model on Geo-SC zhang2023telling , reusing all default hyperparameters that come with the official implementation [^1] https://github.com/Junyi42/geoaware-sc, e.g., training for $2$ epochs using AdamW loshchilov2017decoupled optimizer with $1.25\times 10^{-3}$ initial learning rate and $1.0\times 10^{-3}$ weight decay, coupled with one-cycle learning rate scheduler smith2019super , with a batch size of $1$. Every $5,000$ iterations, models are evaluated on the validation split, and the best performing model is retained. For evaluation, unless stated otherwise, a soft-argmax window of size 15 is used.
Geo-SC specific hyperparameters are also untouched, with the contrastive $\mathcal{L}_{\text{sparse}}$ and dense $\mathcal{L}_{\text{dense}}$ objectives with gaussian noise being used, as well as feature maps dropout and pose-variant augmentation. We refer to the original publication and official implementation for in-depth description of these features.
Our complete loss term is $\mathcal{L}_{\text{sparse}}+\mathcal{L}_{\text{dense}}+\mathcal{L}_{\mathcal{P}}+0.3\times\mathcal{L}_{\mathcal{Z}}+\mathcal{L}_{\text{geom}}$, where $\mathcal{L}_{\mathcal{P}}$, $\mathcal{L}_{\mathcal{Z}}$, and $\mathcal{L}_{\text{geom}}$ correspond to Eq. 3, Eq. 4 and Eq. 6 respectively. The justification for setting the weight $\lambda_{\mathcal{Z}}=0.3$ is provided in Appendix B.
@@ Appendix B Ablations
Table A1: Average PCK@0.1 on SPair-71k validation set for different ablations.
@@
We perform ablations of our designs in Table A1, and report results on the the SPair-71k min2019spair validation set which helps us chose the best performing model. It is not possible to ablate individual loss terms as they each have a distinct purpose without which the prototype cannot properly be learned: $\mathcal{L}_{\mathcal{P}}$ optimizes $\mathcal{P}$, $\mathcal{L}_{\mathcal{Z}}$ optimizes $\mathcal{Z}$, and $\mathcal{L}_{\text{geom}}$ provides a dense supervision signal, i.e., a loss for $\Phi(I,\bm{u})$ when $\bm{u}$ is an arbitrary object pixel, i.e., not a keypoint.
@@ C.1 Additional Metrics
KAP mariotti2024improving PCK only considers matches when both ground-truth points are visible and does not penalize systems that predict strong similarities for points that do not correspond, for instance between the two opposite sides of a car. KAP reformulate the correspondence evaluation as a binary classification problem between the pixels that are close to the target and those those that are not. Crucially, it penalizes high predictions when a source keypoint is invisible in the target.
@@
Table A2: Evaluation under robust metrics. All metrics use per image averaging, and all models use window soft-argmax. All models are trained on SPair-71k, and models with a double dagger benefit from AP-10K pretraining. Models in the $\mathcal{K}$ category use keypoint supervision, while $\mathcal{K}$ do not. Best results are bolded and second best are underlined.
@@
Results in Table A2 confirm the patterns observed in Section 5.2. For all metrics, supervised models performances drop back down to unsupervised-level or worse when evaluated outside their training labels. Interestingly, KAP scores do not widely vary between supervised and unsupervised models, indicating that supervised models are still likely to predict strong similarity between points when none exists.
@@ C.2 Additional Visualizations
We visualize more canonical surfaces in Fig. A1. While the shapes are sensible, we observe some limitations in adequately modeling categories with extreme deformations like birds: points belonging to the wings are predicted close to the body when they are folded, and away when they are spread. However, this is consistent with SPair-71k labeling, where the tips are only labeled when the wings are spread.
We also show some predictions for the unsupervised DINOv2+SD, Geo-SC, and our model on SPair-U in Fig. A3. We observe some interesting failure cases: on the aeroplane, the unsupervised model correctly matches the door, while both supervised models incorrectly predict a training keypoint. In two occasions, Geo-SC predicts points outside of the object when queried on points that are far from training annotations (cow and person). Finally, two very challenging cases are shown with the chair and the tv, illustrating that generic semantic correspondence is still a particularly challenging task.
@@ Appendix D Additional SPair-U Details
We annotated images using the VGG Image Annotator dutta2019vgg . We further post-processed the annotations into JSON files replicating the structure of SPair-71k annotations, i.e., per-image annotations and a list of testing pairs. This allows SPair-U to function as a drop-in replacement for SPair-71k evaluation in any semantic correspondence evaluation script. Note that it is designed to be a benchmark of unseen semantic points intended for evaluating the generalization ability of SC models, therefore does not come with a training or validation split. We present the full list of keypoint semantics of SPair-U in Table A3, per-category statistics in Table A4, and some keypoint visualization in Fig. A2.
@@
[…]

### v2

@@ Appendix A Additional Implementation Details
General implementation. We base our model on Geo-SC zhang2023telling, reusing all default hyperparameters that come with the official implementation [^1] https://github.com/Junyi42/geoaware-sc, e.g., training for $2$ epochs using AdamW loshchilov2017decoupled optimizer with $1.25\times 10^{-3}$ initial learning rate and $1.0\times 10^{-3}$ weight decay, coupled with one-cycle learning rate scheduler smith2019super, with a batch size of $1$. Every $5,000$ iterations, models are evaluated on the validation split, and the best performing model is retained. For evaluation, unless stated otherwise, a soft-argmax window of size 15 is used.
@@
Point cloud sampling. When backprojecting image points using Eq.˜2, we obtain a point cloud whose size depends on the number of visible object pixels. This can cause an imbalance of the samples in the final loss terms, with larger objects contributing more. Furthermore, these point clouds have a very specific grid-like structure inherited from the bitmap format of images, which is dense on surfaces parallel to the image plane and gets sparser as the angle increases. Therefore, we first subsample each training point cloud to a size of $k=1024$ points using farthest point sampling to obtain a fixed number of well-distributed samples.
When computing the local geometry loss $\mathcal{L}_{\text{geom}}$, we use neighborhoods of size $k^{\prime}=64$. An ablation study of this parameter is provided in Table˜1(c). We also provide in Algorithm˜1 the pseudocode for the sampling strategy used to obtain $\mathbf{B}_{r}$ in the posed space.
Input: Point cloud $\mathcal{X}$, seed point $p$, number of neighbors $k$
Output: neighbor set $B$
// Start with the seed
$B\leftarrow$ {pp};
while $|B|<k$ do
 // Filter out already selected points
 $\mathcal{X}^{\prime}\leftarrow\mathcal{X}\setminus B$;
 for $x$ in $\mathcal{X}$ do
    // Compute the distance to the closest point in $B$
    $D_{x}\leftarrow\min_{y\in B}\left\lVert x-y\right\rVert_{2}$;
   end for
 // Add the point with minimal distance to $B$
 $B\leftarrow B\cup\{\arg\min_{x\in\mathcal{X}^{\prime}}D_{x}$};
end while
return $B$;
Algorithm 1 Pseudo-geodesic sampling
@@ Appendix B Ablations
Table A1: Results of different ablations.
(a) Average PCK@0.1 on SPair-71k test set and SPair-U for different models.
DINOv2+SD Geo-SC losses Canonical prototype Spair-71k Spair-U
76.5 60.0
85.6 57.1
75.9 66.0
85.4 66.1
(b) Average PCK@0.1 on SPair-71k validation set for general ablations.
Ablation PCK
@@
(c) Average PCK@0.1 on SPair-71k validation set for different neighborhood size.
Neighborhood size PCK
4 85.5
8 86.1
16 85.9
32 86.5
64 86.5
128 86.6
256 86.1
(d) Average PCK@0.1 on SPair-71k validation set for different rigidity constraints.
Selected points PCK
3 86.6
4 86.4
5 86.1
6 86.4
7 86.5
8 86.0
all 86.5
General ablations. We perform ablations of our designs in Table˜1(d), and report results on the the SPair-71k min2019spair validation set which helps us chose the best performing model. It is not possible to ablate individual loss terms as they each have a distinct purpose without which the prototype cannot properly be learned: $\mathcal{L}_{\mathcal{P}}$ optimizes $\mathcal{P}$, $\mathcal{L}_{\mathcal{Z}}$ optimizes $\mathcal{Z}$, and $\mathcal{L}_{\text{geom}}$ provides a dense supervision signal, i.e., a loss for $\Phi(I,\bm{u})$ when $\bm{u}$ is an arbitrary object pixel, i.e., not a keypoint. We can however examine the different effect of our and Geo-SC’s specific loss terms on both SPair-71k and SPair-U, by comparing a simple supervised approach using a DINOv2 and SD backbone trained only using $\mathcal{L}_{\text{sparse}}$, adding only the Geo-SC specific losses, adding only our canonical prototype losses, and our proposed model that combines them. Results shown in Table˜1(a) show a clear pattern, i.e., adding our canonical prototype loss results in a very small drop in performance on seen keypoints (i.e., SPair-71k), which we attribute to the models having less capacity to fully overfit the training keypoint supervision, but endows them with the ability to generalize much better to unseen points (i.e., SPair-U). Conversely, the Geo-SC losses allow models to perform really well on seen keypoints but its effect on generalization ranges from harmful to null (with vs. without Geo-SC losses). These results also demonstrate that our contributions do not require Geo-SC to work, as they also boost performance of the supervised baseline on unseen keypoints (with vs. without Canonical prototype).
@@
Neighborhood size in $\mathbf{C}_{q}$ and $\mathbf{B}_{r}$. We experiment with different neighborhood sizes when computing $\mathcal{L}_{\text{geom}}$ and valite the different models on the SPair-71k validation set in Table˜1(c). Results show little to no effect of the neighborhood size, which is consistent with our previous finding that our losses do not improve performance on semantics points that are present in the training set.
Number of points in Eq.˜3 To compute $\mathcal{L}_{\mathcal{P}}$, we compute a global rigid transformation between the posed and canonical keypoints, which our qualitative analysis in Fig.˜6 and Fig.˜A1 shows to be reasonable despite being a coarse simplification of the problem. In order to evaluate its impact quantitatively, we evaluate altered version of this procedure where we learn $\mathcal{P}$ not by globally aligning all keypoints but only a randomly sampled local subset of them. Results in Table˜1(d) show that only considering a local neighborhood does not impact the validation performance of our model.
@@ C.1 Additional Metrics
KAP mariotti2024improving PCK only considers matches when both ground-truth points are visible and does not penalize systems that predict strong similarities for points that do not correspond, for instance between the two opposite sides of a car. KAP reformulates the correspondence evaluation as a binary classification problem between the pixels that are close to the target and those those that are not. Crucially, it penalizes high predictions when a source keypoint is invisible in the target.
@@
Table A2: Evaluation under robust metrics. All metrics use per-image averaging, and all models use window soft-argmax. All models are trained on SPair-71k, and models with a double dagger benefit from AP-10K pretraining. Models in the $\mathcal{K}$ category use keypoint supervision, while $\mathcal{K}$ do not. Best results are bolded and second best are underlined.
@@
Results in Table˜A2 confirm the patterns observed in Section˜5.2. For all metrics, supervised models performances drop back down to unsupervised-level or worse when evaluated outside their training labels. Interestingly, KAP scores do not widely vary between supervised and unsupervised models, indicating that supervised models are still likely to predict strong similarity between points when none exists.
@@ C.2 Additional Visualizations
We visualize more canonical surfaces in Fig.˜A1. While the shapes are sensible, we observe some limitations in adequately modeling categories with extreme deformations like birds: points belonging to the wings are predicted close to the body when they are folded, and away when they are spread. However, this is consistent with SPair-71k labeling, where the tips are only labeled when the wings are spread.
We also show some predictions for the unsupervised DINOv2+SD, Geo-SC, and our model on SPair-U in Fig.˜A3. We observe some interesting failure cases: on the aeroplane, the unsupervised model correctly matches the door, while both supervised models incorrectly predict a training keypoint. In two occasions, Geo-SC predicts points outside of the object when queried on points that are far from training annotations (cow and person). Finally, two very challenging cases are shown with the chair and the tv, illustrating that generic semantic correspondence is still a particularly challenging task.
@@ Appendix D Additional SPair-U Details
We annotated images using the VGG Image Annotator dutta2019vgg. We further post-processed the annotations into JSON files replicating the structure of SPair-71k annotations, i.e., per-image annotations and a list of testing pairs. This allows SPair-U to function as a drop-in replacement for SPair-71k evaluation in any semantic correspondence evaluation script. Note that it is designed to be a benchmark of unseen semantic points intended for evaluating the generalization ability of SC models, therefore does not come with a training or validation split. We present the full list of keypoint semantics of SPair-U in Table˜A3, per-category statistics in Table˜A4, and some keypoint visualization in Fig.˜A2.
@@
M v [Uncaptioned image] Ω T [Uncaptioned image] [Uncaptioned image] Avg
@@
DINOv2+SD Geo-SC Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure A4: Visualization matches for randomly selected object points. Colors are provided as a way to distinguish the points.